# Anti-Phishing Training (Still) Does Not Work: A Large-Scale Reproduction of Phishing Training Inefficacy Grounded in the NIST Phish Scale

### v2

@@ 3.1 Study Design
Figure 2: Illustration of our methodology. Our two-factor design randomly assigns different trainings to different subjects, and then randomly sends phishing lures of varying complexity to the subjects. We compared between-subject performance using a mix of standard and novel metrics to test our five hypotheses.
@@
Our study design is illustrated in Figure˜2. We employed a two-factor between-subjects experimental design [47]. Subjects were randomly assigned to one of three experimental conditions using our organization’s training management system. Each condition reflected the phishing training they received. Subjects were randomly selected, trained, and later received simulated phishing emails of varying complexity. We quantitatively compared subjects’ performance in handling these phishing emails between the treatments.
@@
Factor 1—Training: The research team collaborated with a cybersecurity training vendor who produces comprehensive security awareness training materials, phishing simulation exercises designed to augment those materials, and implementation assistance for organizational deployment. This partnership enabled access to professional training content and established phishing templates.
@@
The vendor offers two distinct training approaches for phishing. The first approach, Treatment A, consists of lecture-based videos on the dangers of various social engineering attacks, including phishing recognition tactics, social engineering techniques, and organizational security protocols. Figure˜3 gives a screenshot. Comprehension is assessed through a quiz. The second approach, Treatment B, consists of those videos followed by a series of interactive phishing exercises that present users with samples from a library of approximately 20 phishing templates. These exercises guide users through identifying suspicious elements and security deficiencies in emails that roughly correspond to the cues assessed by the NIST Phish Scale, and reinforces appropriate response behaviors such as the “Report Phish” button in the email client [19]. Finally, we included a Control Group whose members received no cybersecurity training prior to the phishing simulation phase. This group was included to establish baseline susceptibility measurements. [^3] Control group participants received the same training materials after completion of the study to ensure organizational compliance with all cybersecurity awareness requirements. The study timing dictated a small control group, as ACME Corp. required that the majority of employees receive annual cybersecurity training for compliance purposes.
Figure 3: Training for Treatment A: Videos and Quiz
@@
Factor 2—Phishing Lure Complexity: After training, all subjects received one simulated phishing email, of varying complexity. To establish controlled and consistent evaluation of participant responses, we systematically assessed and categorized the difficulty of each simulated phishing email using a structured expert review process informed by the NIST Phish Scale framework [27].
@@
Phishing templates were selected from the vendor’s library. We evaluated these templates using the official NIST Phish Scale User Guide methodology to ensure accurate difficulty classification. For the Phishing Cues dimension, raters assessed observable indicators that might alert users to fraudulent content, including spelling errors, suspicious URLs, formatting inconsistencies, and sender authenticity markers [16]. For the Premise Alignment dimension, raters evaluated how well each email’s subject matter, sender, and content aligned with typical organizational communications that ACME Corp. employees receive in their regular workflow [2].
@@
The assessment team comprised three members of ACME Corp.’s IT security staff, each experienced in identifying real-world phishing attacks. Following Downs et al.’s advice, two team members independently evaluated each phishing email template [30]. Any disagreements regarding difficulty classification were discussed and resolved, incorporating a third rater as needed. This collaborative validation process ensured that our phishing simulation represented a controlled range of difficulty levels essential for evaluating training effectiveness across standardized complexity measures [8].
@@
Our final simulation campaign used nineteen phishing templates that we divided into Easy (high cues/low alignment — 5721 emails), Medium (some cues, some alignment — 2279 emails) and Hard (few cues/high alignment — 4511 emails). Examples are given in Figure˜4. A full summary is given in Table˜6 in the Appendix.
(a) Easy: low premise, high cues.
(b) Hard: high premise, low cues.
Figure 4: Two Example Phishing Templates.
@@
3.2 Subjects
Study Context: The subjects were employees at a U.S.-based international financial technology (“fintech”) firm. Participants represented a diverse cross-section of the organization, including employees from finance, technology, operations, customer service, and administrative functions. All participants were full-time employees with active email accounts who regularly use organizational communication systems.
We excluded participants who: (1) had inactive email accounts during the simulation period, (2) were on extended leave during the study period. Additionally, participants who received phishing simulations but had technical delivery failures (bounced emails, system errors) were excluded.
Subject Recruitment: This study leveraged organizational security training data collected as part of routine cybersecurity awareness training. The training was introduced through official organizational communications emphasizing its importance for maintaining security compliance and protecting organizational assets. This training was mandatory due to the organization’s need for regulatory compliance (cf. Appendix˜2.1.3).
@@ 3.2 Subjects
Randomization: Subjects were randomly assigned to each treatment. To mitigate bias by organizational role or function, we used stratified sampling [91] across departments to ensure representative distributions in each treatment group.
Completion Rates: Our data came from the 12,511 participants who completed the full experimental protocol. Table˜3 shows the completion rates by treatment type. The comparable completion rates across conditions (ranging from 87.3% to 89.1%) suggest that our random sample was preserved between recruitment and completion.
Table 3: Recruitment pool and completion rates by condition. Note the comparable drop-out rates by condition.
@@
Table 4: Summary of study hypotheses, their theoretical grounding, and relation to prior work. The hypotheses address the effects of phishing email difficulty, training, and training modality on detection and reporting behaviors.
@@
H1 (Phish Scale Effect): Phishing emails with higher difficulty, as measured by the NIST Phish Scale, will have higher click-through rates than less difficult ones. Grounded in the foundational NIST Phish Scale framework [27] and validation studies [8]. Supported by phishing cue research [16] and systematic email analysis [99]. The NIST Phish Scale has been validated in lab settings [8]. We provide the first large-scale enterprise validation.
H2 (Training Impact–General): Training will increase the detection and reporting of phishing emails. Evidence of training efficacy, reduced click rates [19], improved detection accuracy [90], and short-term benefits [56]. Disputed [48]. Tests whether training effects from small-scale studies [19, 90] replicate at scale, addressing Franz et al.’s call [37].
H3 (Interactive Training Effect–Reporting): Interactive components yield higher reporting rates than lecture-only. From Franz et al.’s modality taxonomy [37], and evidence of the effectiveness of active and feedback-rich training [19, 56, 82]. Builds on academic findings of interactive training benefits [19, 56]. We provide the first test at organizational scale.
H4 (Interaction Effect): There will be an interaction between training modality and phishing difficulty, with interactive training more beneficial for easier lures. Emerges from cognitive load and decision strategy research [30], difficulty-aware detection patterns [99], and findings on human limitations [2]. Addresses the gap identified by Franz et al. [37] regarding interaction effects between training and phishing complexity, largely unexplored empirically.
H5 (Reporting Timeliness): Training will improve the timeliness of phishing reporting, with trained users reporting suspicious emails more quickly than untrained users. Grounded in studies showing that training enhances threat recognition speed [19]. Supported by evidence that security education improves decision-making efficiency [90]. While individual reporting speed has been measured [53], no studies have examined training effects on organizational-level reporting timeliness [88].
3.3 Hypotheses
Based on prior work on phishing training efficacy [19, 18, 44, 90, 37, 48] and documented differences in risk perception [2, 29], we formulated and tested five hypotheses. Table˜4 states each hypothesis, its basis, and its relation to prior work.
Our specific hypotheses are:
  1. 1.
@@ 3.3 Hypotheses
With respect to hypothesis novelty and prior testing: H1 has been tested in small academic settings [8] but never at enterprise scale. H2 and H3 build on established training effectiveness research [19, 90, 56] but applies these findings in a large-scale operational context using a standardized difficulty measure; H4 is novel, addressing the research gap identified by Franz et al. [37] regarding interactions between training modality and phishing complexity, which has received minimal empirical attention in prior work. H5 is also novel, directly addressing Steves et al.’s research question about organizational temporal patterns [88].
3.4 Metrics
Our outcome measures were a mix of standard metrics for comparison to prior results [19, 18], and novel metrics related to response timeliness per Steves et al. [88]. We performed measurements using the vendor’s simulation platform, which reported the standard metrics directly and timing information that we used to derive the novel metrics.
Standard Metrics: We employed three standard behavioral metrics to assess training effectiveness [19, 18]:
  • Open Rate is the percent of recipients who open the simulated phishing email, measured via embedded tracking pixels. [^4] This is the standard definition, though we acknowledge it can be influenced by email client settings and security configurations [6].
  • Click-Through Rate (CTR) is the percent of recipients who click on malicious links within the phishing emails. CTR is our primary indicator of phishing susceptibility [90].
  • Reporting Rate is the percent of recipients who report the suspicious email to the organization’s security team using the integrated “Report Phish” button, representing proactive security behavior and successful threat recognition [29].
Novel Temporal Metrics: Abroshan et al. suggested examining temporal patterns in user responses [2], and Steves et al. asked: “Is time to first report sooner than time to first click?” [88]. Rapid reporting of phish attempts enables the feedback loop built into modern email defenses (Figure˜1). We introduce the following metrics to see organizational-level temporal security dynamics:
@@ 3.4 Metrics
  • Campaign Inoculation Status determines whether an organization can mitigate a phishing campaign before exploitation occurs:
@@
  • Organizational Inoculation Index quantifies the temporal advantage (or disadvantage) of reporting over clicking:
@@
    Positive values indicate “inoculation” where reports preceded clicks, enabling organizational threat mitigation before exploitation. Negative values indicate that clicks occurred before reports, suggesting vulnerability windows.
@@
This novel organizational-level metric directly addresses Steves et al.’s research question while providing actionable insights into collective security behaviors that complement individual-focused assessments. The Organizational Inoculation Index enables organizations to evaluate whether their workforce’s collective reporting behavior provides sufficient early warning to enable proactive threat response.
3.5 Analysis
@@
Data Validation and Quality Assurance: Prior to analysis, we conducted data validation to ensure analytical integrity. This included verification of random assignment effectiveness across groups, assessment of technical delivery success rates, and examination of potential outliers or data anomalies. Participants with incomplete data due to technical failures, extended leave, or incomplete training were excluded from analysis to maintain dataset quality [8].
@@
Statistical Approach: We employed two complementary analytical approaches to examine training effectiveness. Our primary analysis used two-way analysis of variance (ANOVA) to test main effects and interactions between training modality (Treatments A, B, and Control) and phishing difficulty level (Easy, Medium, Hard) on our outcome measures. Note that although ANOVA with binary outcomes violates standard normality assumptions, this approach aligns with established practices in phishing research [19, 18] and enables direct comparison with prior studies.
@@
To address the limitations of ANOVA with binary data, we conducted confirmatory analyses using logistic regression models appropriate for binary outcomes (clicked/not, reported/not) [83]. The logistic regression models employed maximum likelihood estimation and provided odds ratios for effect size interpretation. Both analytical approaches yielded consistent conclusions regarding training effectiveness, with ANOVA providing F-statistics for comparison with prior literature and logistic regression offering more statistically appropriate modeling of our binary outcomes. We report both approaches to provide comprehensive statistical evidence while acknowledging the methodological trade-offs inherent in each approach. All analyses were conducted using Python [77] with pandas [57] for data manipulation, NumPy [45] for numerical computations, SciPy [95] for statistical tests, and statsmodels [83] for regression modeling and maximum likelihood estimation. Visualizations were created using matplotlib [50] and seaborn [97].
3.6 Threats to Validity
@@
We encourage the reader to consider several limitations when interpreting the results of our study. We distinguish these limitations into three kinds of threats to validity. Following the guidance of Verdecchia et al., we focus on what we perceive to be the most crucial ones [94].
@@
Construct Validity: Construct validity is concerned with whether we operationalized the desired constructs of our study appropriately [84, 98]. In this study, we worked with two primary constructs: (1) the particular interpretations of the two different types of phishing training we deployed, which depend on our cybersecurity training vendor; and (2) our assessment of the difficulty of phishing lures.
It is possible that our cybersecurity vendor’s training was simply lackluster, hence its failure to produce any significant improvement in subject performance. However, note that our results align with prior work that used different trainings. Meanwhile, to control for subjectivity in rating phishing lures, we applied the NIST Phish Scale and used multiple independent analysts to judge difficulty [85].
@@
Internal Validity: Internal validity threats affect the reliability of conclusions we have about the relationship between the cause and the effect we observe [84, 98]. For our study, we used stratified random sampling to mitigate potential biases related to job role and other organizational issues. Due to the large sample sizes for Treatments A and B, our study has significant statistical power. Organizational constraints required us to limit the sample size for our control group. This group was still large enough to provide a baseline.
@@
Like all real-world phishing measurement studies, our work faces several measurement concerns. False positive clicks may occur when automated security systems (security gateways, antivirus, etc.), shadow IT applications, or personally installed tools / browser extensions automatically scan or prefetch URLs from emails, generating click events without actual user interaction. False negative reporting may also occur, wherein users report phishing emails through alternative channels, such as contacting the IT helpdesk directly, rather than the simulation platform’s designated “Report Phish” button. These behaviors would, respectively, incorrectly increase and decrease our measurements of subject interactions. We cannot fully eliminate such threats, but our large sample size mitigates their impact on our overall conclusions.
@@
External Validity: External threats impact the generalizability of our research [81, 100]. Our study was situated in the USA at a fintech company. The fintech industry context may limit generalizability to other sectors, as financial services employees may have different baseline levels of security awareness when compared to other industries. However, given the minimal effect sizes measured in our study, it is unclear how large of a threat this is. The study’s focus on immediate post-training effects also limits our understanding of long-term training impact. Future research should examine whether the modest effects we observed persist over time or diminish as training recency decreases, or if they improve with volume or other dimensions.
4 Results

### v3

@@ 3. Methodology
Figure 1. Illustration of our methodology. Our two-factor design randomly assigns different trainings to different subjects, and then randomly sends phishing lures of varying complexity to the subjects. We compared between-subject performance using a mix of standard and novel metrics to test our five hypotheses. :
To summarize the preceding analysis: phishing training is widely practiced, though recent evidence sheds doubt on its utility. Few studies have rigorously assessed outcomes against standardized difficulty measures at sufficient scale for robust inference. We aim to reproduce, testing hypotheses derived from prior works.
3.1. Study Design
@@
Our study design is illustrated in Figure 1. We employed a two-factor between-subjects experimental design (hillmanEvaluatingOrganizationalPhishing2023, ). Subjects were randomly assigned to one of three experimental conditions using our organization’s training management system. Each condition reflected the phishing training they received. Subjects were randomly selected, trained, and later received simulated phishing emails of varying complexity. We quantitatively compared subjects’ performance in handling these phishing emails between the treatments.
@@
Factor 1: Training: The research team collaborated with a cybersecurity training vendor who produces comprehensive security awareness training materials, phishing simulation exercises designed to augment those materials, and implementation assistance for organizational deployment. This partnership enabled access to professional training content and established phishing templates.
@@
The vendor offers two distinct training approaches for phishing. The first approach, Treatment A, consists of lecture-based videos on the dangers of various social engineering attacks, including phishing recognition tactics, social engineering techniques, and organizational security protocols. Comprehension is assessed through a quiz. The second approach, Treatment B, consists of those videos followed by a series of interactive phishing exercises that present users with samples from a library of approximately 20 phishing templates. These exercises guide users through identifying suspicious elements and security deficiencies in emails that roughly correspond to the cues assessed by the NIST Phish Scale, and reinforces appropriate response behaviors such as the “Report Phish” button in the email client (carellaImpactSecurityAwareness2017, ). Finally, we included a Control Group whose members received no cybersecurity training prior to the phishing simulation phase. This group was included to establish baseline susceptibility measurements. [^3] Control group participants received the same training materials after completion of the study to ensure organizational compliance with cybersecurity awareness requirements. The study timing dictated a small control group, as ACME Corp. required that most employees receive annual cybersecurity training for compliance purposes.
@@
Factor 2: Phishing Lure Complexity: After training, all subjects received one simulated phishing email, of varying complexity. To establish controlled and consistent evaluation of participant responses, we systematically assessed and categorized the difficulty of each simulated phishing email using a structured expert review process informed by the NIST Phish Scale framework (dawkinsScalingNISTPhish2023, ).
@@
Phishing templates were selected from the vendor’s library. We evaluated these templates using the official NIST Phish Scale User Guide methodology to ensure accurate difficulty classification. For the Phishing Cues dimension, raters assessed observable indicators that might alert users to fraudulent content, including spelling errors, suspicious URLs, formatting inconsistencies, and sender authenticity markers (canhamNotAllVictims2024, ). For the Premise Alignment dimension, raters evaluated how well each email’s subject matter, sender, and content aligned with typical organizational communications that ACME Corp. employees receive in their regular workflow (abroshanPhishingHappensTechnology2021, ).
@@
The assessment team comprised three members of ACME Corp.’s IT security staff, each experienced in identifying real-world phishing attacks. Following Downs et al.’s advice, two team members independently evaluated each phishing email template (downsDecision2006, ). Any disagreements regarding difficulty classification were discussed and resolved, incorporating a third rater as needed. This collaborative validation process ensured that our phishing simulation represented a controlled range of difficulty levels essential for evaluating training effectiveness across standardized complexity measures (barrientosScalingPhishAdvancing2021, ).
@@
Our final simulation campaign used nineteen phishing templates that we divided into Easy (high cues/low alignment — 5721 emails), Medium (some cues, some alignment — 2279 emails) and Hard (few cues/high alignment — 4511 emails). A full summary is given in Table 6 in Appendix C.
@@
3.2. Subjects
Study Context: The subjects were full-time employees at a U.S.-based international financial technology (“fintech”) firm. Subjects represented a diverse cross-section of the organization, including employees from finance, technology, operations, customer service, and administrative functions.
Subject Recruitment: This study leveraged organizational security training data collected as part of routine cybersecurity awareness training. The training was introduced in official organizational communications emphasizing its importance for maintaining security compliance and protecting organizational assets. This training was mandatory due to the organization’s regulatory environment.
@@ 3.2. Subjects
Randomization: Subjects were randomly assigned to each treatment. To mitigate bias by organizational role or function, we used stratified sampling (sutterAvoidingHookInfluential2022, ) across departments to ensure representative distributions in each treatment group.
Completion Rates: Our data came from the 12,511 participants who completed the full experimental protocol. Table 2 shows the completion rates by treatment type. Comparable completion rates across conditions (ranging from 87.3% to 89.1%) suggest our random sample was preserved between recruitment and completion.
Table 2. Recruitment pool and completion rates by condition.
@@
Table 3. Summary of study hypotheses, their theoretical grounding, and relation to prior work. The hypotheses address the effects of phishing email difficulty, training, and training modality on detection and reporting behaviors.
@@
H1–Phish Scale Effect: Phishing emails with higher difficulty, as measured by the NIST Phish Scale, will have higher click-through rates than less difficult ones. Grounded in the foundational NIST Phish Scale framework (dawkinsScalingNISTPhish2023, ) and validation studies (barrientosScalingPhishAdvancing2021, ). Supported by phishing cue research (canhamNotAllVictims2024, ) and systematic email analysis (wrightPhishingSusceptibilityContext2023, ). The NIST Phish Scale has been validated in lab settings (barrientosScalingPhishAdvancing2021, ). We provide the first large-scale enterprise validation.
H2–Training Impact (General): Training will increase the detection and reporting of phishing emails. Evidence of training efficacy, reduced click rates (carellaImpactSecurityAwareness2017, ), improved detection accuracy (sumnerExaminingFactorsImpacting2022, ), and short-term benefits (marshallExploringEvidenceEmail2024, ). Disputed (hoUnderstandingEfficacyPhishing2025, ). Tests whether training effects from small-scale studies (carellaImpactSecurityAwareness2017, ; sumnerExaminingFactorsImpacting2022, ) replicate at scale, addressing Franz et al.’s call (franzSoK, ).
H3–Interactive Training Effect (Reporting): Interactive components yield higher reporting rates. Per Franz et al.’s analysis (franzSoK, ), and evidence of effectiveness of active and feedback-rich training (carellaImpactSecurityAwareness2017, ; marshallExploringEvidenceEmail2024, ; schoniYouKnowWhat2024, ). Lab findings show interactive training benefits (carellaImpactSecurityAwareness2017, ; marshallExploringEvidenceEmail2024, ). We provide the first test at organizational scale.
H4–Interaction Effect: There will be an interaction between training modality and phishing difficulty, with interactive training more beneficial for easier lures. Emerges from cognitive load and decision strategy research (downsDecision2006, ), difficulty-aware detection patterns (wrightPhishingSusceptibilityContext2023, ), and findings on human limitations (abroshanPhishingHappensTechnology2021, ). Addresses the gap identified by Franz et al. (franzSoK, ) regarding interaction effects between training and phishing complexity, largely unexplored empirically.
H5–Reporting Timeliness: Trained users will report suspicious emails more quickly than untrained users. Grounded in studies showing that training enhances threat recognition speed (carellaImpactSecurityAwareness2017, ). Supported by evidence that security education improves decision-making efficiency (sumnerExaminingFactorsImpacting2022, ). While individual reporting speed has been measured (lain2022phishing, ), no studies have examined training effects on organizational-level reporting timeliness (stevesCategorizingHumanPhishing2020, ).
3.3. Hypotheses
Based on prior work on phishing training efficacy (carellaImpactSecurityAwareness2017, ; caputoGoingSpearPhishing2014, ; haneyComplianceImpactTracing2024, ; sumnerExaminingFactorsImpacting2022, ; franzSoK, ; hoUnderstandingEfficacyPhishing2025, ) and documented differences in risk perception (abroshanPhishingHappensTechnology2021, ; doingAnalysisPhishingReporting2024, ), we formulated and tested five hypotheses. Table 3 states each hypothesis, its basis, and its relation to prior work. We hypothesize that:
  1. (1)
@@ 3.3. Hypotheses
With respect to hypothesis novelty and prior testing: H1 has been tested in small academic settings (barrientosScalingPhishAdvancing2021, ) but never at enterprise scale. H2 and H3 build on established training effectiveness research (carellaImpactSecurityAwareness2017, ; sumnerExaminingFactorsImpacting2022, ; marshallExploringEvidenceEmail2024, ) but applies these findings in a large-scale operational context using a standardized difficulty measure; H4 is novel, addressing the research gap identified by Franz et al. (franzSoK, ) regarding interactions between training modality and phishing complexity, which has received minimal empirical attention in prior work. H5 is also novel, directly addressing Steves et al.’s research question about organizational temporal patterns (stevesCategorizingHumanPhishing2020, ).
3.4. Metrics
Our outcome measures were a mix of standard metrics for comparison to prior results (carellaImpactSecurityAwareness2017, ; caputoGoingSpearPhishing2014, ; lopezaguilarPhishingVulnerabilityPersonality2025, ) and novel metrics related to response timeliness per Steves et al. (stevesCategorizingHumanPhishing2020, ). We performed measurements using the vendor’s simulation platform, which reported the standard metrics directly and timing information that we used to derive the novel metrics.
Standard Metrics: We employed three standard behavioral metrics to assess training effectiveness (carellaImpactSecurityAwareness2017, ; caputoGoingSpearPhishing2014, ):
  • Open Rate is the percent of recipients who open the simulated phishing email, measured via embedded tracking pixels. [^4] This is the standard definition, though we acknowledge it can be influenced by email client settings and security configurations (antoniouSystematicMethodExecute2015, ).
  • Click-Through Rate (CTR) is the percent of recipients who click on malicious links within the phishing emails. CTR is our primary indicator of phishing susceptibility (sumnerExaminingFactorsImpacting2022, ).
  • Reporting Rate is the percent of recipients who report the suspicious email to the organization’s security team using the integrated “Report Phish” button, representing proactive security behavior and successful threat recognition (doingAnalysisPhishingReporting2024, ).
Novel Temporal Metrics: Abroshan et al. suggested examining temporal patterns in user responses (abroshanPhishingHappensTechnology2021, ), and Steves et al. asked: “Is time to first report sooner than time to first click?” (stevesCategorizingHumanPhishing2020, ). Rapid reporting of phish attempts enables the feedback loop built into modern email defenses (Figure 3). We introduce the following metrics to see organizational-level temporal security dynamics:
@@ 3.4. Metrics
  • Campaign Inoculation Status determines if an organization can mitigate a phishing campaign before exploitation occurs:
@@
  • Organizational Inoculation Index (OII) quantifies the temporal advantage (or disadvantage) of reporting over clicking:
@@
    Positive values indicate “inoculation” where reports preceded clicks, enabling organizational threat mitigation before exploitation. Negative values instead suggest vulnerability windows.
@@
This novel organizational-level metric directly addresses Steves et al.’s (stevesCategorizingHumanPhishing2020, ) research question while providing actionable insights into collective security behaviors that complement individual-focused assessments. The Organizational Inoculation Index lets organizations evaluate whether their workforce’s collective reporting behavior provides early warning to enable proactive threat response.
3.5. Analysis
@@
Data Validation and Quality Assurance: Prior to analysis, we conducted data validation to ensure analytical integrity. This included verifying random assignment effectiveness across groups, assessing technical delivery success rates, and examining potential outliers or data anomalies (barrientosScalingPhishAdvancing2021, ).
@@
Statistical Approach: We employed two complementary analytical approaches to examine training effectiveness. Our primary analysis used two-way analysis of variance (ANOVA) to test main effects and interactions between training modality (Treatments A, B, and Control) and phishing difficulty level (Easy, Medium, Hard) on our outcome measures. Note that although ANOVA with binary outcomes violates standard normality assumptions, this approach aligns with established practices in phishing research (carellaImpactSecurityAwareness2017, ; caputoGoingSpearPhishing2014, ) and enables direct comparison with prior studies.
@@
To address the limitations of ANOVA with binary data, we conducted confirmatory analyses using logistic regression models appropriate for binary outcomes (clicked/not, reported/not) (seabold2010statsmodels, ). The logistic regression models employed maximum likelihood estimation and provided odds ratios for effect size interpretation. Both analytical approaches yielded consistent conclusions regarding training effectiveness, with ANOVA providing F-statistics for comparison with prior literature and logistic regression offering more statistically appropriate modeling of our binary outcomes. We report both approaches, while acknowledging the methodological trade-offs inherent in each approach.
3.6. Threats to Validity
@@
We encourage the reader to consider several limitations when interpreting the results of our study (verdecchia2023threats, ). We distinguish these limitations into three kinds of threats to validity.
@@
Construct Validity: Construct validity is concerned with whether we operationalized the desired constructs of our study appropriately (shadishCookCampbell2002, ; wohlinEtAl2012, ). In this study, we worked with two primary constructs: (1) the phishing training we deployed, which depends on our cybersecurity training vendor; and (2) our assessment of the difficulty of phishing lures. First, though vendors may vary in training design, our cybersecurity vendor’s training is regulation-compliant, and our results align with prior work that used different trainings. Second, to control for subjectivity in rating phishing lures, we used the NIST Phish Scale with multiple analysts to judge difficulty (sjobergBergersen2023, ).
@@
Internal Validity: Internal validity threats affect the reliability of conclusions we have about the relationship between the cause and the effect we observe (shadishCookCampbell2002, ; wohlinEtAl2012, ). For our study, we used stratified random sampling to mitigate potential biases related to job role and other organizational issues. Due to the large sample sizes for Treatments A and B, our study has significant statistical power. Organizational constraints required us to limit the sample size for our control group. This group was still large enough to be a baseline.
@@
Like all real-world phishing measurement studies, our work faces several measurement concerns. False positive clicks may occur when automated security systems (security gateways, antivirus, etc.), “shadow IT” applications, or personally installed tools / browser extensions automatically scan or prefetch URLs from emails, generating click events without actual user interaction. False negative reporting may also occur, in which users report phishing emails through alternative channels, such as contacting the IT help-desk directly, rather than the simulation platform’s designated “Report Phish” button. These behaviors would, respectively, incorrectly increase and decrease our measurements of subject interactions. We cannot fully eliminate such threats, but our large sample size mitigates their impact on our overall conclusions.
Although such measurement noise is inherent in operational studies, the large sample size and balanced distribution across conditions mitigate its systematic influence on our results. Every effort was made to standardize user behavior, including explicit instruction during both training modalities on how to identify and report suspicious messages through the organization’s approved “Report Phish” mechanism. Future work might explicitly quantify these instrumentation effects by correlating simulation telemetry with downstream SOC logs or ticketing data to better understand their prevalence and direction of bias.
@@
External Validity: External threats impact the generalizability of our research (runesonHost2009, ; yin2018, ). Our study was situated in the USA at a fintech company. The context of the fintech industry may limit generalizability to other sectors, as financial services employees may have a higher baseline level of security awareness. However, given the minimal effect sizes measured in our study, this threat is not substantial. Our focus on immediate post-training effects also limits our understanding of long-term training impact. Future work might assess the effect of training recency and volume.
3.7. Ethics
We acknowledge the importance of informed consent as part of an empirical study. Employees are informed at onboarding that periodic phishing simulations are a mandatory condition of employment, consistent with industry practice in financial services organizations (doingAnalysisPhishingReporting2024, ). Our study then analyzed data collected as part of routine, required security awareness training, using the existing corporate platform and controls.
Our protocol was reviewed by both the company and the partner academic institution’s IRB. Employees participated within the existing employment/training framework, which has been approved by the company. In consultation with management, we excluded high-distress scenarios (e.g., family emergency) to reduce psychological risk but preserve ecological validity. Because the research used de-identified operational data generated by routine commercial operations, our institutional IRB concluded that from the academic partner’s view our study did not constitute human subjects research.
4. Results